# Learning to Infer Parameterized Representations of Plants from 3D Scans

### v1

@@ 6.3 Skeletonization
Synthetic Noise Depth images
[53] $0.561\pm 0.312$ $0.701\pm 0.344$ $0.595\pm 0.324$
[6] $0.544\pm 0.286$ $0.583\pm 0.301$ $0.716\pm 0.377$
[30] $0.740\pm 0.355$ $0.776\pm 0.363$
Ours $\mathbf{0.037\pm 0.052}$ $\mathbf{0.034\pm 0.044}$ $\mathbf{0.042\pm 0.055}$
Table 3: Comparison to Xu et al. [53], Chaudhury et al. [6], and Livny et al. [30] for 3D skeleton extraction in terms of Chamfer Distance. “-” means that the method crashed due to numerical problems.
We quantitatively evaluate our method for 3D skeleton extraction from input point cloud using the bidirectional Chamfer Distance. Chamfer Distance helps capturing the geometrical discrepancy between the ground truth skeleton and the predicted skeleton by measuring the sum of nearest neighbour distances in both directions.
We compare our method with two classical plant skeletonization baselines Livny et al.[30], Xu et al. [53] and a stochastic skeleton refinement method Chaudhury et al. [6] that takes a predicted skeleton as input and outputs a refined skeleton. In this comparison we apply the skeleton refinement on the output skeletons from applying Xu et al. [53] on the plant point clouds.
Figure 7: Qualitative results for skeletonization by our method, Livny et al. [30], Xu et al. [53], and Chaudhury et al. [6] refinement on Xu skeleton, from clean synthetic, noisy and partial point clouds taken from monocular depth maps. “-” means that the method crashed due to numerical problems.
All these methods are designed to take scans of leafless plants as input. However, our dataset contains Chenopodium Album plants with leaves. Table 3 shows the results on the test sets of synthetic clean, noisy and partial point clouds. Our method shows consistent performance on the different test sets, suggesting robustness to noise and missing parts in the input point clouds for the skeletonization task. Since the baseline methods are not designed to take leaves as input, they tend to interpret leaf geometry as part of the skeleton, while the ground truth skeleton mainly describes the skeletal structure of the stem, branches and petioles. This explains the higher Chamfer distances, and the visual difference shown in Fig. 7.
@@ 6.4 Segmentation
Synthetic
PlantNet 91 97 94 88 98 99 99 97 89 76 82 70 97 94 95 91
PSegNet 91 99.5 95 90 99 99.6 99 98 95 88 92 85 99 89 94 88
Ours 89 93 91 84 97 98 98 95 68 54 59 45 97 98 97 94
Noise
PlantNet 88 81 84 73 93 99 96 91 75 32 45 29 74 89 80 68
PSegNet 99 75 85 74 90 99.7 95 90 76 21 33 20 98 80 88 79
Ours 85 90 88 78 97 97 97 94 60 47 51 37 93 89 91 84
@@
Table 4: Comparison to PlantNet [26] and PSegNet [25] for semantic segmentation on synthetic and noisy points test sets using standard classification measures Precision, Recall, F1-Score and Intersection over Union (IoU). All values are percentages.
We quantitatively evaluate our method’s performance on the semantic segmentation task applied to our test sets using standard classification metrics: precision, recall, F1 score, and Intersection over Union (IoU). For each class, precision is defined as the ratio of true positives to the sum of true positives and false positives, while recall is the ratio of true positives to the sum of true positives and false negatives. The F1 score is the harmonic mean of precision and recall. Finally, IoU is computed as the ratio of the intersection to the union of the predicted and ground truth point sets for a given class.
We compare our method to the two strong baselines PlantNet [26] and PSegNet [25]. Both PlantNet and PSegNet take a deep neural network approach with special sampling strategies, and are suitable for semantic and instance segmentation of plants.
Table 4 shows the results on the test sets of synthetic and noisy point clouds. We do not test segmentation methods on the depth image test set, as we do not have ground truth labels in this case. For synthetic scans, PSegNet outperforms all the methods except on Cotyledons segmentation, where our method shows a better performance. However, when noise is introduced, our method outperforms the other methods for most of the measures. Note that the petiole is the most difficult part to segment; this is because of its cylindrical shape, which is very close to the stem branches and makes it difficult to discriminate between these two classes. Overall, our method is robust to noise and performs on-par with the strong baselines.
Fig. 8 shows the semantic segmentation outputs of our method compared to the ground truth, where the semantics are properly captured even for noisy inputs. Furthermore, we show our method’s ability to perform instance segmentation in Fig. 9, where it successfully separates individual plant parts and remains robust to noise.
Figure 8: Examples of semantic segmentation results on synthetic clean and noisy point clouds.
Figure 9: Examples of instance segmentation results on synthetic clean and noisy point clouds.
7 Conclusion and Discussion
In this paper we have presented a data-driven method to infer plants parametric representations i.e. L-Strings, from 3D unstructured point cloud input. We have shown that the L-String representation contains structural and geometric information that the input scans lack, and that such information allows for multiple tasks like 3D reconstruction, skeletonization and segmentation. Our results have shown that our method is on-par with the state-of-the-art for these tasks on synthetic data, and is robust to noisy and incomplete input. However, the performance of our method is limited to synthetic data, and first results from this first version of the method need to be qualitatively and quantitatively improved when applied to real scans.
Figure 10: Our 3D reconstruction results on input scans of real Chenopodium Album plants.
This work demonstrate the feasibility to learn latent space of 3D plant structures and reconstruct plant architecture parametric models from it. In the next step, our aim is to upscale our method to address fully real 3D plant point clouds. For this we need to investigate the level of accuracy required to simulated plants real plants so that the method can use simulated plants to learn real plant architectures. Fig 10 shows first tests of our 3D reconstructions from real scans. Although the reconstructions are not yet accurate on all levels, global information is learned by our model.
8 Acknowledgements
We thank Franck Hétroy-Wheeler for helpful discussions, and Ayan Chaudhury, Frédéric Boudon, and Aymen Merrouche for providing support for their code. This work was partially supported by French government funding managed by the National Research Agency under grant ANR-24-CE23-1586 (4DPlants).

### v2

@@ 6.2 Skeletonization
Full Branch
Method Clean Noisy Depth images Clean
[10.1145/1289603.1289610] $0.0102$ $0.0110$ $0.0145$
[Chaudhury2020.02.15.950519] $0.0139$ $0.0154$ $0.0449$
[livny2010automatic] $0.0257$ $0.0282$ -
Ours $0.0178$ $0.0174$ $0.0199$ $0.0161$
Table 3: Comparison for skeletonization w.r.t. Chamfer Distance. “-”: method crashed due to numerical problems. ✗: method’s scale cannot operate on the branch level.
We quantitatively evaluate 3D skeleton extraction using the commonly used bidirectional Chamfer Distance. We compare our method to two classical plant skeletonization baselines Livny et al. [livny2010automatic], Xu et al[10.1145/1289603.1289610] and a stochastic skeleton refinement method Chaudhury et al[Chaudhury2020.02.15.950519] that takes a predicted skeleton as input and outputs a refined skeleton. The skeleton refinement is applied on the output skeletons of Xu et al[10.1145/1289603.1289610]. All these methods are designed to take scans of leafless plants as input, and their output cannot be adjusted to different scales. In contrast, our method can output skeletons with or without the main veins of the leaves.
Figure 7: Comparison to Xu et al[10.1145/1289603.1289610] for skeletonization on noisy input. Ours is applicable at different scales, and can output a full or a branch skeleton, while achieving visually accurate results.
Table 3 shows the results for both the full skeleton that includes leaf veins and the branch skeleton that only includes the branching structure. Our method shows consistent performance on the different test sets, suggesting robustness to noise and missing parts in the input point clouds. Since baseline methods are not designed for input with leaves, they cannot output skeletons at different scales. Fig. 7 shows a visual comparison on noisy input with the best method from table 3 Xu et al[10.1145/1289603.1289610], where our method achieves results close to the ground truth. More visual comparisons are in the appendix. Fig. 6 (last column) shows skeletons extracted from scans of real plants. The skeletons closely follow the topology in the scans.
@@ 6.3 Segmentation
We compare our method to the strong baselines PlantNet [LI2022243] and PSegNet [psegnet2022]. PSegNet provides labels on sparsely sampled point clouds, which we transfer to the full input point cloud using nearest neighbors similarly to what we do for our approach. Fig. 8 shows the semantic segmentation results on examples from all test sets, where each type of plant organ is assigned a unique color. Our method is robust to noise and partial data, is on-par with the baselines, and outperforms PSegNet on the petioles segmentation. Quantitative comparisons on semantic segmentation and qualitative results on instance segmentation can be found in appendix. Fig. 6 (third column) shows the segmentation of scans of real plants with our method. Note that all organs are well segmented overall.
Figure 8: Comparison to PlantNet [LI2022243] and PSegNet [psegnet2022] for semantic segmentation on clean, noisy and partial point clouds.
7 Conclusion
In this paper we have presented a data-driven method to infer parametric representations of plants i.e. L-Strings, from 3D unstructured point cloud input. This advancement was possible by leveraging procedural models both for designing the neural networks to learn a parametric shape space for 3D plants, and for generating synthetic training data. We have shown that the L-String representation contains structural and geometric information that the input scans lack, and that such information allows for multiple tasks like 3D reconstruction, skeletonization and segmentation at once, with performance on-par with strong baselines for each individual task. Our results have further shown that our method, trained purely on synthetic data, generalizes well to noisy scans of real plants.
8 Acknowledgements
We thank Franck Hétroy-Wheeler for helpful discussions, and Ayan Chaudhury and Frédéric Boudon for sharing their codes. This work was partially supported by French government funding managed by the National Research Agency under grant ANR-24-CE23-1586 (4DPlants).
Appendix
9 Implementation Details
We implemented our method using Pytorch on Quadro RTX 5000 GPU. We learn a latent space $\mathcal{S}$ of dimension $dim_{\mathcal{S}}=64$ with the recursive encoder-decoder pair architecture in Fig. 4 in the main paper, of hidden layer size $h_{recur}=128$, and the classifiers with a hidden layer $h_{class}=128$. We optimize the weights of the network using Adam optimizer with a learning rate of 0.001, decaying by factor of 0.5 every 50 epochs. For the point cloud encoder we train a PointNet point regression network with Adam optimiser on a learning rate of 0.001. The weights of the layers of all the trained networks are initialized using Xavier uniform initialization.
Our method works on binary trees, while the L-String trees in our dataset do not have a binary tree structure in general. We summarize the L-Strings into binary trees, by combining the modules that always occur together as individual nodes with concatenated parameter sets. In particular, we combine the two cotyledon modules into one Cotyledons node with 6 parameters. Stems are always followed by a petiole and a leaf, therefore we combine them in one node called Stem that has 13 parameters. In case of a branching, the branch module is combined with the stem, petiole and leaf modules to form a Branch node of 15 parameters. Finally, the first stem module of the tree forms a node on its own called Root of 5 parameters. Parameters that are drawn from a Gaussian distribution do not correlate with their parent in the tree structure and cannot be learned by the recursive network, thus they are set to the mean during the training and are then optimized with other parameters in the test-time optimization phase.
For the segmentation experiments where we use the $k$ nearest neighbors algorithm, we set $k=10$ for our results, and $k=3$ for PSegNet results. As for the reconstruction experiments, when using SIREN we need to input the point normals along with the point cloud, for this we use a normal estimation PCA-based method, which fits local planes to each point’s neighborhood.
10 Additional evaluations
10.1 Semantic Segmentation
We quantitatively evaluate our method’s performance on the semantic segmentation task applied to our test sets using standard classification metrics precision, recall, F1 score, and Intersection over Union (IoU). For each class, precision is defined as the ratio of true positives to the sum of true positives and false positives, while recall is the ratio of true positives to the sum of true positives and false negatives. The F1 score is the harmonic mean of precision and recall. Finally, IoU is computed as the ratio of the intersection to the union of the predicted and ground truth point sets for a given class. Table 4 shows the results on all test sets. Our method is robust to noise and performs overall on-par with the strong baselines PlantNet [LI2022243] and PSegNet [psegnet2022] on each individual task, while performing all the tasks at once.
@@ 10.1 Semantic Segmentation
Clean
PlantNet 85.3 88.2 86.7 76.6 98.4 99.5 98.9 97.9 82.1 71.7 76.5 62.0 98.9 94.4 96.7 93.5
PSegNet 99.1 99.3 99.2 98.5 98.4 98.9 98.6 97.3 66.6 59.6 62.9 45.9 99.7 97.0 98.6 97.3
Ours 84.8 88.0 86.2 75.9 98.0 98.4 98.2 96.5 66.7 59.4 62.2 45.9 96.7 95.8 96.2 92.7
Noisy
PlantNet 84.2 88.4 86.2 75.8 97.7 99.1 98.4 96.9 77.6 62 69.0 52.6 96.4 93.1 94.7 89.9
PSegNet 99.3 83.6 90.7 83.1 95.9 99.9 97.9 95.9 96.2 29.6 45.3 29.3 99.6 98.1 98.9 97.7
Ours 83.0 86.3 84.5 73.3 97.9 98.3 98.1 96.2 64.2 57.4 59.9 43.7 95.8 93.9 94.8 90.2
Depth maps
PlantNet 86.2 86.8 86.5 76.2 98.5 99.2 98.8 97.7 70.2 65.6 67.8 51.3 97.4 88.7 92.9 86.7
PSegNet 99.1 98.9 99.0 98.1 96.2 99.9 98.1 96.2 98.9 54.9 70.6 54.6 99.8 98.7 99.2 98.5
Ours 84.6 84.2 84.0 73.2 98.4 98.6 98.5 97.0 57.9 54.1 54.5 38.8 96.8 94.8 95.7 92.0
@@
Table 4: Comparison to PlantNet [LI2022243] and PSegNet [psegnet2022] for semantic segmentation on clean and noisy points test sets using standard classification measures Precision, Recall, F1-Score and Intersection over Union (IoU). All values are percentages.
10.2 Instance Segmentation
We qualitatively compare our method to PlantNet [LI2022243] and PSegNet [psegnet2022] for the Instance segmentation task. Fig. 9 shows the instance segmentation results on examples from all test sets, where each individual plant organ is assigned a unique color. Note that the colors are to distinguish the segmented organs without any correspondence between the results and the ground truth. Our method is on-par with the baselines for this task where it is able to segment individual organ instances and is robust to noise and partial data.
Figure 9: Visual comparison between our method, PlantNet and PSegNet for the instance segmentation task. Different colors are assigned to segmented instances without correspondence to the ground truth.
10.3 Skeletonization
To support the results shown in Table 3 in the main paper on the skeletonization task, we qualitatively compare to the baselines Livny et al. [livny2010automatic], Xu et al. [10.1145/1289603.1289610] and Chaudhury et al. [Chaudhury2020.02.15.950519] on all test sets. Fig. 11 shows the visual qualitative results of all methods on extracting a full skeleton from an input point cloud, with our results of extracting the branch skeleton compared to the ground truth. Our results are close to the ground truth skeletons and are robust to noise and partial input on both scales.
10.4 Influence of Latent Dimension
All loss functions are necessary by construction for our models to train and thus cannot be ablated. Instead, we study the influence of the dimension of the learned latent space $\mathcal{S}$ on the performance of our model. For that, we evaluate the L-string latent representation stage by training the recursive auto-encoders with different latent dimensions $dim_{\mathcal{S}}$ and measuring the Tree-Edit Distance [ferraro-hal-00827474] between the input L-strings and the reconstructions. Tree-Edit distance measures the difference between two tree graphs w.r.t. the number of operations needed to translate from one tree structure to the other, with additional local cost functions that is the difference between the node parameters in our case, hence it measures the difference in both topology and shape between plants represented as L-strings. Table 5 shows the average reconstruction error in tree-edit distance for different dimensions, in our implementation we choose $dim_{S}$ that is most compact with minimum error.
$dim_{\mathcal{S}}$ Tree Edit Distance
32 0.227
64 0.196
128 0.197
256 0.296
Table 5: Analysis of the influence of latent dimension on the performance of the L-string auto-encoder model w.r.t. the Tree-Edit Distance.
10.5 Latent Space Analysis
To analyse the data distribution in the learned latent space $\mathcal{S}$, we project the latent points representing plants from the training set in Fig. 10 using UMAP [mcinnes2020umapuniformmanifoldapproximation], the uniform manifold approximation and projection non-linear technique for dimension reduction with local and global structure preservation. Different plant structures are assigned different colors in Fig. 10, one can notice that our model learns to encode plants sharing the same tree structure close by in $\mathcal{S}$.
Figure 10: Analysis of the distribution of different tree structures from the training data in the latent space projected in 2D using UMAP. Each color represent a unique plant structure in the training data. Plants sharing the same structure form clusters in $\mathcal{S}$.
Figure 11: Comparison to Livny et al[livny2010automatic], Xu et al[10.1145/1289603.1289610], and Chaudhury et al[Chaudhury2020.02.15.950519] refinement on Xu skeleton for skeletonization. “-” means that the method crashed due to numerical problems. Note that ours is the only method applicable at different scales, and able to output a full skeleton or a branch skeleton, while achieving visually accurate results.